OpenStudy (anonymous):

Is there any relation between ln(-1) and sqr root(-1)

hartnn (hartnn):

ln(-1) = i\(\pi \)

hartnn (hartnn):

from e^{i\(\pi\)} = -1

OpenStudy (anonymous):

Oh...... ya

OpenStudy (anonymous):

Let i^i = e^x -> i ln i = x i ln (sqrt(-1)) = x i/2 ln -1 = x but since ln -1 = ipi x = -pi/2

OpenStudy (anonymous):

WELL THANX GUYS

OpenStudy (anonymous):

So, 

OpenStudy (anonymous):

That is also TRUE?

hartnn (hartnn):

yep.

OpenStudy (anonymous):

So, all numbers do have relation?

hartnn (hartnn):

euler's relation is great!

OpenStudy (anonymous):

agreed.